# Enhancing Visual Reasoning with Autonomous Imagination in Multimodal Large Language Models

### v1

@@ 5.4 Object Placement
As shown in Tab. 1, our method consistently outperforms baseline methods in both object locating and placement performance. It is important to note that our cursor-only ablation baseline is identical to our full method when locating an object. Therefore, we assign them with the same locating result. Additionally, the sampling method is intractable for locating, we directly utilize the locating results of our method for placement.
@@
Methods lacking advanced visual reasoning capabilities perform poorly on the placement task, as the successful placement requires inferencing about the correct spatial position for that element beyond visual recognition. The cursor-only baseline, which moves visual marker instead of scene modification, does not perform as effectively as our complete method. This difference underscores the advantage of scene manipulation in simplifying the reasoning process by revealing additional visual information. Although the sampling baseline receives substantial visual information, it still under-performs relative to our method, and even falls short of the cursor-only baseline. Conversely, in simpler jigsaw puzzle solving, the sampling method achieves a relatively high success rate compared to the cursor-only baseline. This contrast highlights the importance of a structured reasoning pathway, particularly in complex, open-world scenarios where the abundant visual information could overwhelm the MLLM, preventing it from identifying the correct answer despite its presence in the sample. By following the reasoning process of closed-loop control, MLLMs can progressively approach the correct answer without requiring an exhaustive number of samples, resulting not only in greater efficiency but also in improved accuracy.
6 Limitations and Future Work
While our method enables MLLMs to perform visual reasoning tasks by operating within a structured framework autonomously, its expressiveness is constrained by the defined set of operators. Future work could focus on extending the operator set to address a broader range of tasks as MLLMs’ capability to handle complex operations improves. Furthermore, if future MLLMs natively possess robust and precise image editing capabilities, the reasoning process could become more efficient and streamlined, potentially removing the need for a specially designed imagination space.
@@ 7 Conclusion
Current MLLMs still struggle with visual reasoning tasks beyond clue finding that humans perform intuitively. We propose a new reasoning paradigm to conduct closed-loop decision-making and reasoning with imagination, thus creating an interactive imagination space where MLLMs can refine reasoning. Our method employs plug-and-play imagination space and operator design, enabling MLLMs to modify visual content autonomously, enabling general and natural CoT construction. We propose the IVRB benchmark across three challenging visual reasoning tasks, simple jigsaw puzzle solving, dense counting, and object placement, demonstrating the limitations of existing methods for reasoning beyond clue finding. Our approach, however, enables MLLMs to handle these tasks with autonomous imagination from their native reasoning abilities, pushing the boundaries of visual reasoning capabilities for MLLMs.
8 Acknowledgement
This work is supported by National Natural Science Foundation of China (U23A20311,62206245).
@@ Appendix A More Experiments
ROPE Benchmark. We evaluate our method using the recently proposed ROPE benchmark for Multi-Object Hallucination [6], focusing on the most challenging subset where GPT-4o exhibits the poorest performance, which is answering about single object given multi-object image, heterogeneous object types, on unseen data split. Our method uses GPT-4o as the base model with our imagination space and operations equipped. Please refer to the original paper for more details. Given that this benchmark emphasizes the identification of abstract objects that are not suitable for segmentation, we turn the focus operation into drawing a large rectangular region by MLLMs natively, through selecting top-left and bottom-right corners. The benchmark indicates that MLLMs are prone to hallucinations when confronted with multiple objects that introduce additional visual distractions. We demonstrate that using our method, MLLMs can autonomously focus on the important region despite these distractions, leading to more accurate responses due to the elimination of irrelevant visual information.
Models Acc. Models Acc.
LLaVA-34B 30.81% GroundHOG* 38.13%
IDEFICS 6.50% Qwen VL-C 15.37%
Yi-VL-34B 0.41% MiniCPM-V 14.39%
CogVLM-C 13.50% GPT-4o 53.74%
GLaMM* 52.28% GPT-4o+Ours 65.00%
Table 2: Comparison of our method with baselines in ROPE benchmark, see Sec. A for details. Except for our method, the results of other baselines are cited from the latest leaderboard in https://multi-object-hallucination.github.io/ before our arXiv submission time. For each base model family, we take the one with the highest performance. Mechanistically grounded LVLMs (marked with *) take visual prompts by dedicated pointer tokens. Please refer to the original paper [6] for more details.
@@
Imagine with Image Editing Models. We experimented with utilizing image editing models that accept natural language input, such as the state-of-the-art model SmartEdit [14]. Ideally, such models would provide effective guidance by generating the target goal based on natural language descriptions. However, as shown in Fig. 7, we find that the image editing model is not ideal when it comes to following complex instructions. This underscores the significant challenge of “imagining” the target state in open-world problems, suggesting that a world model capable of accurately providing such visual guidance requires further development.
@@
Please also refer to the following figures for additional qualitative experiment results. We present qualitative results on tasks including dense counting, simple jigsaw puzzle solving (please also see the video in project page), and object placement. These images more effectively illustrate the detailed visual reasoning processes and the changes within the imagination space.
(a) SmartEdit placing pear
(b) SmartEdit placing apple
Figure 7: We show that existing image editing models lack the capability to effectively comprehend and execute complex commands, such as placing the apple/pear at the left/right side on the table.
Figure 8: We show failure cases in our method caused by hallucination. These include instances where MLLMs mistakenly perceive a fry that does not exist, misinterpret the wrongly focused floor as a dumbbell, and incorrectly believe that a puzzle piece has been placed in the correct spot.
Appendix B More Experiment Details
B.1 Dense Counting Experiments in Fig. 5(a) and 5(b)
The aim of these experiments was to evaluate the impact of difficulty and reasoning steps on performance. To achieve this, we selected a subset of cases from dense counting and supplemented them with simpler cases involving fewer objects. We then re-conducted the experiments on this partial dataset using GPT-4o, alongside a newly introduced textual CoT reasoning baseline, as discussed in the main paper.
In the first experiment (Fig. 5(a)), we show the success rate of different methods when faced with varying numbers of objects. In the second experiment (Fig. 5(b)), we test how different methods improve when given different limits on the number of reasoning steps. The result is shown as the solvable rate, which indicates the percentage of cases the method can solve under the current budget. When imposing such a limit, our model is restricted to performing no more than the specified number of steps, with each step defined as a combination of the focus operation and the subsequent operation performed after focusing. For the textual CoT baseline, we explicitly instructed the model on the maximum number of reasoning steps it was allowed to use.
B.2 Simple Jigsaw Puzzles
The rotations of pieces are not considered in the task. As multiple trials are allowed, we restrict the total number of attempts to 20 and report the finish rate, defined as the proportion of missing pieces successfully placed. The sampling methods used involve comparing pairs of target locations and iteratively selecting the better option until a single target location remains. We guarantee that the correct answer exists within the sampling process. In our method, once a jigsaw piece is moved to the imagined target location, it is considered placed. The GPT-4o baseline selects the target location by outputting coordinates, while the VCoT baseline determines the target location by drawing a block, leveraging its specialized training for such tasks. For baselines that directly output a location, such as GPT-4o and VCoT, we ensure that the jigsaw puzzle piece is provided in the given image. Our method locates and selects a piece independently, as it possesses this capability, while other methods are automatically given a piece.
B.3 Object Placement
Similar to the approach used in solving simple jigsaw puzzles, the methods compared include the following: GPT-4o, which outputs the target location by specifying coordinates; the sampling method, which iteratively compares sampled results until a single option remains; VCoT, which determines placement by drawing blocks; and our method, which outputs the final placement after completing a transformation operation. In alignment with how related works in robotics handle sampling-based methods [18, 8], the sampling baseline restricts its search to a sub-region of the space, filtering out certain incorrect answers. For example, regions occupied by existing objects are excluded from consideration. This principle is also incorporated into the transformation operation in our method. Specifically, a focus operation is first applied to identify the platform where the target object should reside, creating a region mask. The transformation operation then ensures that no movement is suggested if it would lead the object outside the defined mask. This approach enhances efficiency and ensures a fair comparison between the sampling baseline and our method by reducing unnecessary exploration of invalid regions.
@@ B.3 Object Placement
Figure 10: Qualitative demonstration of the counting process autonomously performed by MLLMs.
Figure 11: Qualitative demonstration of the counting process autonomously performed by MLLMs.

### v3

@@ 4.5 Object Placement
Similar to the approach used in solving simple jigsaw puzzles, the methods compared include the following: GPT-4o and o1, which output the target location by specifying coordinates; the sampling method, which iteratively compares sampled results until a single option remains; VCoT and Molmo, which determines placement by drawing blocks or markers; and our method, which outputs the final placement after completing a transformation operation. In line with how related work in robotics handles sampling-based methods (Kapelyukh et al., 2024; Ding et al., 2024), the sampling baseline restricts its search to a subregion of the space, filtering out certain incorrect answers. For example, regions occupied by existing objects are excluded from consideration. This principle is also incorporated into the transformation operation in our method. Specifically, a focus operation is first applied to identify the platform on which the target object should reside, creating a region mask. The transformation operation then ensures that no movement is suggested if it would lead the object outside the defined mask. This approach improves efficiency and ensures a fair comparison between the sampling baseline and our method by reducing unnecessary exploration of invalid regions.
@@
As shown in Tab. 1, our method consistently outperforms the baseline methods in both object locating and placement performance. It is important to note that our cursor-only ablation baseline is identical to our full method when locating an object. Therefore, we assign them with the same locating result. Additionally, the sampling method is intractable for locating. Thus, we directly utilize the locating results of our method for placement. Methods lacking advanced visual reasoning capabilities perform poorly on the placement task, as successful placement requires inferencing about the correct spatial position for that element beyond visual recognition. The cursor-only baseline, which moves the visual marker instead of the scene modification, does not perform as effectively as our complete method. This difference underscores the necessity for substantial modification in visual task decomposition. Although the sampling baseline receives substantial visual information, it still underperforms relative to our method and even falls short of the cursor-only baseline. In contrast, in simple jigsaw puzzle solving, the sampling method achieves a relatively high success rate compared to the cursor-only baseline. This contrast highlights the importance of a structured reasoning pathway, particularly in complex, open-world scenarios where the abundant visual information could overwhelm the MLLM, preventing it from identifying the correct answer despite its presence in the sample. By following the closed-loop control reasoning process, MLLMs can progressively approach the correct answer without requiring an exhaustive number of samples, resulting in greater efficiency and improved accuracy.
4.6 Multi-Object Hallucination Benchmark
We evaluated our method using the recently proposed ROPE benchmark for Multi-Object Hallucination (Chen et al., 2024), focusing on the most challenging subset where GPT-4o exhibits the poorest performance, which is answering about a single object given multi-object image, heterogeneous object types, on unseen data split. Our method uses GPT-4o as the base model that equips the imagination space. Please refer to the original paper for more details. Given that this benchmark emphasizes the identification of abstract objects that are not suitable for segmentation, we turn the focus operation into drawing a large rectangular region by MLLMs natively, through selecting top-left and bottom-right corners. The benchmark indicates that MLLMs are prone to hallucinations when confronted with multiple objects that introduce additional visual distractions. We demonstrate that using our method, MLLMs can autonomously focus on the important region despite these distractions, leading to more accurate responses due to the elimination of irrelevant visual information.
[…]
Models Acc. Models Acc. Models Acc. Models Acc.
LLaVA-34B 30.81% CogVLM-C 13.50% Qwen VL-C 15.37% Molmo 28.90%
IDEFICS 6.50% GLaMM* 52.28% MiniCPM-V 14.39% o1 60.20%
Yi-VL-34B 0.41% GroundHOG* 38.13% GPT-4o 53.74% GPT-4o+Ours 65.00%
Figure 7: Failure cases in our method caused by hallucination, including instances where MLLMs mistakenly perceive a fry that does not exist, misinterpret the wrongly focused floor as a dumbbell, and incorrectly believe that a puzzle piece has been placed in the correct spot.
4.7 Discussions
Failure cases. In Fig. 7, we illustrate failure cases of our approach, which are majorly resulted from the hallucination of MLLMs.
Time cost. The time cost of our approach depends mainly on three factors: the running time of a single call of imagination operators, a single call of MLLM reasoning, and the number of MLLM calls. The calls of imagination operators are quite efficient. For example, each call of the SAM model takes 220ms on NVIDIA RTX 4090 on average, which is very efficient compared with MLLM reasoning. Considering the number of calls for MLLM reasoning, the average numbers of MLLM API call for our approach are: counting=117 (for all objects), jigsaw puzzle=20 (for one piece), object placement=15 (for one object), multi-object hallucination=26 (for one object). As shown by the comparisons on GPT-4o Sampling baseline in Tab. 1, our approach is more efficient than previous sampling-based imagination strategies (Kapelyukh et al., 2024; Ding et al., 2024), the closest previous studies to ours, by enabling to solve broader tasks with vast solution spaces.
@@ 5 Conclusion
In this work, we target at tackling visual reasoning problems using MLLMs where textual-to-visual conversion is the major bottleneck. We propose the autonomous imagination approach, which employs plug-and-play imagination space and operator design, enabling MLLMs to modify visual content autonomously, leading to closed-loop decomposition of this conversion process. We conduct experiments under visual reasoning tasks that are beyond the perceptual capabilities of MLLMs, while remained straightforward to humans. The results show that these tasks can be effectively tackled natively by MLLMs equipped with our approach.
Limitations and future work. Our work is a proof of concept: while closed-loop task decomposition overcomes the visual-to-textual conversion bottleneck in our demonstrated tasks, resource constraints prevented us from training a model with native closed-loop reasoning and visual modification capabilities. This also prevents us from validating our approach under large, general-purpose visual benchmarks. We hope that our work inspires future efforts to close this critical gap.
Acknowledgement
This work was supported by National Natural Science Foundation of China (U23A20311,62206245).
@@ Appendix A More Experiments
Imagine with Image Editing Models. We experimented with utilizing image editing models that accept input of natural languages, such as the state-of-the-art model SmartEdit (Huang et al., 2024). Ideally, such models would provide effective guidance by generating the target goal based on descriptions in natural language. However, as shown in Fig. 8, we find that the image editing model is not ideal when it comes to following complex instructions. This underscores the significant challenge of “imagining” the target state in open-world problems, suggesting that a world model capable of accurately providing such visual guidance requires further development.
CogCoM. We have also experimented with a potentially feasible baseline CogCoM (Qi et al., 2024). However, we find that it does not function well when tackling our challenges. Despite explicitly trained with a GROUNDING ability for counting objects, as shown in Tab. Autonomous Imagination: Closed-Loop Decomposition of Visual-to-Textual Conversion in Visual Reasoning for Multimodal Large Language Models, directly prompting the model to output the counting result is more accurate than performing GROUNDING. When asked to complete the jigsaw puzzle by finding target location, it starts hallucinating by outputing unrelated text, as shown in Fig. 9. This is potentially due to the model trained on different tasks and not generalizing to our challenged tasks. We decided to present the result in the appendix instead.
Appendix B Additional Qualitative Results
@@
We present qualitative results on tasks that include counting, solving simple jigsaw puzzles, and placing objects in Fig. 1014. These images more effectively illustrate the detailed visual reasoning processes and the changes within the imagination space. Please also see the supplementary video, which shows the reasoning process of our approach in the jigsaw puzzle solving task.
Appendix C System Prompts
We provide system prompts for the reasoning instructions given to the MLLMs in our tasks, as shown in Fig. 15, as well as the prompts used as reminders for autonomous imagination operations, as illustrated in Fig. 16.
@@ Appendix C System Prompts
Figure 11: Qualitative demonstration of the counting process autonomously performed by our approach.
Figure 12: Qualitative demonstration of the counting process autonomously performed by our approach.
@@
Figure 15: General reasoning instructions provided to models for visual reasoning, as well as the reasoning instruction given to utilize autonomous imagination. It is not difficult for an advanced MLLM to figure out this plan. As we aim to break the visual-to-textual conversion bottleneck, we simplify this by directly given the reasoning plan to the MLLM to avoid occasional failure caused by the MLLM choosing a bad reasoning strategy.
Figure 16: Detailed prompts instructing the MLLM about the operations it can utilize. This is given to the MLLM as a reminder when the MLLM enters corresponding state during the reasoning.